OpenStudy (anonymous):

y= (5x^2-3)^8 differentiate with respect to x

OpenStudy (anonymous):

\[\frac{\partial \left(5 x^2-3\right)^8}{\partial x}=80 x \left(5 x^2-3\right)^7 \]